Manufacturing Overhead Allocation Rate


Andrews Company manufactures blenders, offering its customers two model options – a basic model and a deluxe model.  Each model of blender goes through the same two-step production process – it is assembled in the Assembly Department and packaged for sale in the Packaging Department.  For the month that is about begin, Andrews Company is estimating manufacturing overhead costs of $680,000 in the Assembly Department and $270,000 in the Packaging Department.  The company also estimates the following usage of direct labor hours and machine hours for the month:

 


Assembly Department

Packaging Department

Direct   Labor Hours

500

2,050

Machine   Hours

1,800

250

 

What is Andrews Company’s plantwide overhead allocation rate if the company uses direct labor hours as its allocation base?

 


 

 

What is Andrews Company’s plantwide overhead allocation rate if the company uses machine hours as its allocation base?



What is the departmental allocation rate for the Assembly Department if the department uses machine hours as its allocation base?

 

 


 

 

What is the departmental allocation rate for the Packaging Department if the department uses direct labor hours as its allocation base?

 

 


 

After the month was over, the company was able to provide you with the following data on actual usage and production:

 


Basic Model

Deluxe Model

Total

Units Produced:

1,000

450

1,450

Assembly   Dept.




     Direct   Labor Hrs.

200

275

475

     Machine   Hrs.

800

1,100

1,900

Packaging   Dept.




     Direct   Labor Hrs.

1,475

675

2,150

     Machine   Hrs.

100

125

225

 

How much manufacturing overhead will be allocated to each model if the company uses a plantwide overhead allocation rate with an allocation base of direct labor hours?

 

 

 

 

 

How much manufacturing overhead will be allocated to each model if the company uses a plantwide overhead allocation rate with an allocation base of machine hours?

 

 

 

 

 

How much manufacturing overhead will be allocated to each model if the company uses departmental overhead allocation rates, with the Assembly Department using an allocation base of machine hours and the Packaging Department using an allocation base of direct labor hours?


 

Andrews Company is considering a change to activity-based costing to account for its manufacturing overhead.  Before the month began, the company’s cost accountants identified the following activities, allocation bases, estimated costs and estimated allocation base usage by model:

 

Activity

Estimated MOH Costs

Allocation Base

Basic Model

Deluxe Model

Total

Setup

$295,000

# of batches

50 batches

15 batches

65 batches

Production

$515,000

Machine hours

700 MHs

900 MHs

1,600 MHs

Testing

$140,000

# of tests

100 tests

30 tests

130 tests

 

Calculate overhead allocation rates for each of the three activities.

-          Setup

 

 

-          Production

 

 

-          Testing

 

 

 

Allocate MOH to each of the models given the following actual data provided by the company:

 

Activity

Basic Model

Deluxe Model

Total

Setup

42 batches

16 batches

58 batches

Production

680 MHs

930 MHs

1,610 MHs

Testing

84 tests

32 tests

116 tests





Units Produced:

1,000 units

450 units

1,450 units

 

MOH Allocated to the Basic Model:

 

 

 

 

 

 

MOH Allocated to the Deluxe Model:
Answer #1

The plantwide allocation rate is calculated by dividing the total overheads of the organization by the total number of hours.

Company's plantwide allocation rate on direct labor hours as the allocation base:

= ($680,000 + $270,000) / (500 + 2,050)
= $950,000 / 2,550
= $372.55 per direct labor hour

Company's plantwide allocation rate on machine hours as the allocation base:

= ($680,000 + $270,000) / (1800 + 250)
= $950,000 / 2,050
= $463.41 per direct labor hour

Departmental allocation rate for Assembly department on basis of machine hours

= $680,000 / 1,800
= $377.78

Departmental allocation rate for Packaging department on basis of direct labor hours

= $270,000 / 2,050
= $131.71

How much manufacturing overhead will be allocated to each model if the company uses a plantwide overhead allocation rate with an allocation base of direct labor hours?

Basic Model:

Total direct labor hours = 200 + 1475 = 1675 hours
Overheads allocated = 1675 hours * $372.55 = $624,021

Deluxe Model:

Total direct labor hours = 275 + 675 = 950 hours|
Overheads allocated = 950 hours * $372.55 = $353,923

How much manufacturing overhead will be allocated to each model if the company uses a plantwide overhead allocation rate with an allocation base of machine-hours?

Basic Model:

Total machine hours = 800 + 100 = 900 hours
Overheads allocated = 900 hours * $463.41 = $417,069

Deluxe Model:

Total machine hours = 1100 + 125 = 1225 hours|
Overheads allocated = 1225 hours * $463.41 = $567,677

How much manufacturing overhead will be allocated to each model if the company uses departmental overhead allocation rates, with the Assembly Department using an allocation base of machine-hours and the Packaging Department using an allocation base of direct labor hours?

Basic Model:

Assembly department:
Overheads allocated = 800 machine hours * $377.78 = $302,224

Packaging department:
Overheads allocated = 1475 direct labor hours * $131.71 = $194,272

Total Overheads = $302,224 + $194,272 = $496,496

Deluxe Model:

Assembly department:
Overheads allocated = 1100 machine hours * $377.78 = $415,558

Packaging department:
Overheads allocated = 675 direct labor hours * $131.71 = $88,904

Total Overheads = $415,558 + $88,904 = $504,462


answered by: DoomWicked